A man has 20 coins consisting of dimes and quarters. If the dimes were quarters and the quarters were dimes, he would have ninety cents more than he has now...how many dimes and quarters does he have?
Answer by checkley75(3666)

D+Q=20
D=20-Q
10D+25Q=10Q+25D+90 NOW SUBSTITUTE (20-Q) FOR D IN THIS EQIUATION
10(20-Q)+25Q=10Q+25(20-Q)+90
200-10Q+25Q=10Q+500-25Q+90
-10Q+25Q-10Q+25Q=500+90-200
30Q=390
Q=390/30
Q=13 QUARTERS.
D+13=20
D=20-13
D=7 DIMES.
PROOF
10*7+25*13=10*13+25*7+90
70+325=130+175+90
395=395